Harry is 4 years older than Marvin. In six years Marvin will be 5/6 as old as Harry. How old is each boy? Now, I think the equation should have been 4 + 5/6 = 6 + 5/6. : Harry is 4 years older than Marvin. In six years Marvin will be 5/6 as old as Harry. How old is each boy? Now, I think the equation should have been 4 + 5/6 = 6 + 5/6.
Answer by nerdybill(743)

Harry is 4 years older than Marvin. In six years Marvin will be 5/6 as old as Harry. How old is each boy?
.
Let M = Marvin's age
then
M+4 = Harry's age
.
then from:
"In six years Marvin will be 5/6 as old as Harry."
we get our formula:
M+6 = (5/6)(M+4+6)
M+6 = (5/6)(M+10)
multiplying both sides by 6:
6M + 36 = 5M + 50
M = 50 - 36
M = 14 (Marvin's age)
M+4 = 18 (Harry's age)